# Privacy-Preserving Spiking Neural Networks: A Deep Dive into Encryption Parameter Optimisation

### v1

@@ 1 Introduction
The human brain, with around 100 billion neurons [22], is the most complex organ, central to sensing, processing, and decision-making. Comprising the cerebrum, cerebellum, and brainstem, it integrates sensory inputs, coordinates bodily functions, and supports higher cognition [22]. This complexity has inspired SNNs, a core paradigm in neuromorphic engineering. Unlike traditional neural networks, SNNs emulate biological neurons by transmitting information as brief electrical spikes [29], [25]. They reproduce synaptic plasticity the foundation of learning and memory and excel at processing spatiotemporal data, enabling adaptive applications such as secure data encryption. Advances in computational neuroscience show how neuronal dynamics emerge from synaptic interactions and spike timing patterns [6], [15]. As the third generation of neural networks, SNNs offer energy efficiency and biologically faithful computation by operating on discrete spikes rather than continuous activations [27], [35]. Their ability to encode both frequency and phase in spike patterns provides a richer representation of data, with applications ranging from visual recognition to probabilistic inference [43]. Despite this promise, adoption has been limited by the lack of scalable and biologically plausible training algorithms. Learning in SNNs is often grounded in Hebbian plasticity, where synaptic strength is modulated by correlations between pre- and post-synaptic activity. The most studied rule, spike-timing-dependent plasticity (STDP), adjusts synaptic weights based on spike timing. Extensions such as triplet and quadruplet STDP incorporate higher-order interactions to capture more complex plasticity. Practical implementations frequently use Poisson spike encoding, where inputs are converted into spike trains proportional to stimulus intensity, though this often results in long encoding times. First-spike encoding provides a faster alternative by leveraging the timing of the earliest spike combined with unsupervised STDP, it enables rapid visual feature extraction [23], [2]. Empirical studies show that such approaches can detect salient features efficiently, while supervised training methods including SpikeProp [4] and its variants—extend backpropagation to spiking dynamics by incorporating leaky membrane potentials and layered architectures. Stabilisation strategies such as low-pass filtering further improve training robustness by reducing noise in error propagation [49]. Together, these developments are advancing the practical deployment of SNNs in neuromorphic systems. Cryptography ensures data integrity, confidentiality, authenticity, and non-repudiation [41]. It transforms raw data into cipher text, which can be decrypted to recover the original information [5]. Broadly, cryptographic methods are categorised into symmetric and asymmetric encryption. Symmetric encryption uses a shared key for both encryption and decryption; it is computationally efficient but dependent on secure key distribution [46]. Asymmetric (or public-key) cryptography employs a key pair (a public key for encryption and a private key for decryption) and supports digital signatures for scalable authentication and key management [48]. Properties and benefits of asymmetric (public-key) cryptography are particularly relevant for SNNs, where protecting data integrity and intellectual property is crucial. Asymmetric cryptography also enables homomorphic encryption, allowing computations directly on encrypted data, which is especially valuable for deploying neural networks securely in cloud or distributed environments [21, 8].
@@ 2 Related works
SNNs stand out from conventional neural networks by processing information through sparse, event-driven spikes rather than continuous activations. Temporal coding enables efficient handling of time-dependent data and reduces energy consumption, as neurons fire only in response to stimuli [15], [16]. The combination of energy efficiency and temporal sensitivity makes SNNs well suited for resource-constrained environments and real-time applications such as robotics, autonomous driving, and decision-making in dynamic contexts. Compared with traditional Artificial Neural Networks (ANNs), which use dense, continuous activations [31], SNNs exploit spike timing and frequency to capture richer spatio-temporal patterns [33]. Such spatio-temporal encoding improves efficiency in domains including vision, speech, and medical diagnostics, while also supporting sustainable, low-power computi ng [52]. The temporal precision inherent in spike-based computation is particularly advantageous for secure data processing, where sparsity and accurate spike timing can enhance encryption and strengthen resilience against attacks.
@@
Synaptic plasticity modulates strength through Long Term Potentiation (LTP) and Long Term Depression (LTD), with transient dynamics captured by Short Term Potentiation (STP) and Short Term Depression (STD). A key mechanism, STDP, adjusts synapses based on pre- and post-synaptic spike timing as illustrated in Figure 3 [2, 47], typically modelled with polynomial or exponential timing-dependent windows. Weight initialisation strongly affects network stability, convergence, and learning performance. Methods include random, interval-based, variance-scaling, data-driven, and hybrid approaches [42]. Balanced initialisation supports STDP, avoiding under- or over-spiking, while lateral inhibition mitigates excessive firing by suppressing neighbouring neurons. Remote Supervised Method (ReSuMe) learning uses an instructive signal to modulate plasticity without altering the post-synaptic membrane [3], ensuring precise spike timing. Effective for both single- and multi-layer networks, it surpasses gradient-based methods like SpikeProp [37] and remains important for supervised SNN tasks.
@@
The Internet of Things (IoT) integrates physical and digital systems, enabling autonomous device communication via Wireless multimedia sensor networks (WMSN), radio frequency identification (RFID), and cloud computing [44, 34]. Wireless sensor networkss (WSNs) provide real-time scalar, image, audio, and video data. Cryptography secures message transmission, converting plaintext to ciphertext with keys and reversing the process on reception [36]. Cryptanalysis identifies vulnerabilities, forming the field of cryptology [11, 24]. Symmetric cryptography employs a single shared key, requiring secret exchange, and offers high speed and low resource use (AES, DES, 3DES, Blowfish). As summarised in Table 1 Asymmetric cryptography uses public-private key pairs, supporting digital signatures and secure key exchange (RSA, Diffie-Hellman, Elliptic Curve Cryptography (ECC), Digital Signature Algorithm (DSA)). Lightweight, bio-inspired, and chaos-based schemes optimise security for constrained devices. AES processes 128-bit blocks with 10, 12, or 14 rounds depending on key size [7], using substitution, shift rows, mix columns, and add round key operations [39, 13]. RSA, introduced in 1977, relies on the difficulty of factorising large numbers and supports one-way encryption for secure transmission and digital signatures [20, 40]. Recent variants enhance wireless sensor network efficiency and biometric encryption, achieving high genuine rates with minimal false positives [9, 38].
@@
A hybrid SNN-RSA approach enhances security and adaptability for data transmission. Input is first encrypted with RSA, processed by an SNN, and transmitted, with decryption reversing the sequence.
3 Methodology
Developing secure and efficient encryption systems is a critical challenge in cybersecurity. This article proposed the use of SNNs for bio-inspired encryption by integrating traditional methods like DES, AES, and RSA to enhance both security and performance. The NEST simulator [17], chosen for its biological accuracy, was used to simulate SNNs, while Python and libraries such as NumPy, Matplotlib, and JSON provided a robust development environment. System configurations included neuron types, simulation time steps, and synaptic plasticity models, with OpenCV used to visualise neural activity and refine the model. Data preprocessing involved converting categorical data into ASCII values suitable for SNN input. For instance, the word ”FIL” was mapped to ASCII values 77, 97, and 110, forming the basis for spike train generation and neuron input currents. Binary representations of ASCII values were transformed into spike trains, where ’1’ indicated a spike and ’0’ indicated no spike (e.g., ASCII 77 = ’01001101’). These spike trains enabled the SNN to process data in a biologically inspired manner. ASCII values were applied as input currents to influence neuronal firing rates and patterns, encoding data into distinct neural activity patterns for different characters and inputs.
@@ 3.3 Network Training and Testing
In the project, employ the ReSuMe algorithm for training the SNN. ReSuMe is particularly well-suited for SNNs because it focuses on adjusting synaptic weights based on the precise timing of spikes, enabling the network to learn temporal patterns with high accuracy. The training involves fine-tuning the learning parameters to achieve optimal performance. For instance, the learning rate, which controls how much the synaptic weights are adjusted in response to each error. Another key consideration during training is the role of STDP, a mechanism where the timing of spikes directly influences synaptic strength. The network’s ability to adjust and refine its response to temporal patterns during training. By monitoring and adjusting STDPs effects, ensure that the network is learning effectively and not just memorising the input patterns. Once training is completed, the network undergoes a rigorous testing phase using a separate validation dataset. It is designed to evaluate how well the network has generalised its learning to new, unseen data. The performance of the SNN is measured using specific metrics such as the Purpura Distance and the Van Rossum metrics.
@@ 3.3.1 Purpura Distance
To compare the dissimilarities between spike trains from different neurons, a method introduces a concept of distance between pairs of spike trains, offering an alternative to the commonly employed rate-based correlation techniques for analysing neuronal responses. Here, spike trains of finite length are represented as points within an abstract space, where a specialised metric assigns a non-negative value $D_{ij}$ to each pair of points $i$ and $j$ [19, 53]. The Victor-Purpura distance possesses key characteristics of a true metric: it is zero only for identical spike trains ($D_{ii}=0$), positive for different spike trains ($D_{ij}>0$ for $i\neq j$), symmetric ($D_{ij}=D_{ji}$), and adheres to the triangle inequality ($D_{ik}\leq D_{ij}+D_{jk}$). The Purpura distance is calculated by determining the minimum cost required to convert one spike train into another through operations such as adding or removing spikes, shifting spike times, or changing the neuron’s identity. The cost associated with each modification is determined by the parameter $q$, which controls the timescale sensitivity for spike shifts [18]. The metric can be generalised into a family of distances, each sensitive to the neuron from which each spike originates. In the analysis, employ the basic Purpura metric with a cost of $q=1$ per unit time for shifting a spike.
@@ 4.1 Benchmark using the cryptography and PyCryptodome libraries
The adoption and development of cryptographic libraries in Python have undergone significant evolution, with libraries like PyCrypto [^1] Available online, https://www.pycrypto.org/, last accessed: 25/08/2025 and Cryptography [^2] Available online, https://github.com/pyca/cryptography, last accessed: 25/08/2025 serving as pivotal milestones. PyCrypto, one of the earliest libraries, provided developers with essential cryptographic tools, including support for symmetric encryption algorithms (eg, AES, DES), hashing algorithms, and public-key cryptography. However, despite its foundational role, PyCrypto suffered from several critical shortcomings. The library’s reliance on insecure defaults, such as AES in Electronic Codebook (ECB) mode, and its auto-generated, poorly structured documentation left developers prone to implementation errors. These gaps not only undermined security but also forced developers to rely on external, often unreliable, sources like blogs and forums, further propagating insecure practices

### v3

@@ 1 Introduction
The human brain, with around 100 billion neurons [25], is the most complex organ, central to sensing, processing, and decision-making. Comprising the cerebrum, cerebellum, and brainstem, it integrates sensory inputs, coordinates bodily functions, and supports higher cognition [25]. This complexity has inspired the development of SNNs, which emulate the brain’s electrical signaling by transmitting information through discrete spike events, forming the foundation of neuromorphic engineering. Unlike traditional neural networks, SNNs emulate biological neurons by transmitting information as brief electrical spikes [32], [28]. They reproduce synaptic plasticity the foundation of learning and memory and excel at processing spatiotemporal data, enabling adaptive applications such as secure data encryption. Advances in computational neuroscience show how neuronal dynamics emerge from synaptic interactions and spike timing patterns [6], [16]. As the third generation of neural networks, SNNs offer energy efficiency and biologically faithful computation by operating on discrete spikes rather than continuous activations [30], [38]. Their ability to encode both frequency and phase in spike patterns provides a richer representation of data, with applications ranging from visual recognition to probabilistic inference [46]. Despite this promise, adoption has been limited by the lack of scalable and biologically plausible training algorithms [24]. Learning in SNNs is often grounded in Hebbian plasticity, where synaptic strength is modulated by correlations between pre- and post-synaptic activity. The most studied rule, spike-timing-dependent plasticity (STDP), adjusts synaptic weights based on spike timing. Extensions such as triplet and quadruplet STDP incorporate higher-order interactions to capture more complex plasticity. Practical implementations frequently use Poisson spike encoding, where inputs are converted into spike trains proportional to stimulus intensity, though this often results in long encoding times. First-spike encoding provides a faster alternative by leveraging the timing of the earliest spike combined with unsupervised STDP, it enables rapid visual feature extraction [26], [2]. Empirical studies show that such approaches can detect salient features efficiently, while supervised training methods including SpikeProp [4] and its variants extend backpropagation to spiking dynamics by incorporating leaky membrane potentials and layered architectures. Stabilisation strategies such as low-pass filtering further improve training robustness by reducing noise in error propagation [53]. Together, these developments are advancing the practical deployment of SNNs in neuromorphic systems. Cryptography ensures data integrity, confidentiality, authenticity, and non-repudiation [44]. It transforms raw data into cipher text, which can be decrypted to recover the original information [5]. Broadly, cryptographic methods are categorised into symmetric and asymmetric encryption. Symmetric encryption uses a shared key for both encryption and decryption; it is computationally efficient but dependent on secure key distribution [50]. Asymmetric (or public-key) cryptography employs a key pair (a public key for encryption and a private key for decryption) and supports digital signatures for scalable authentication and key management [52]. Properties and benefits of asymmetric (public-key) cryptography are particularly relevant for SNNs, where protecting data integrity and intellectual property is crucial. Asymmetric cryptography also enables homomorphic encryption, allowing computations directly on encrypted data, which is especially valuable for deploying neural networks securely in cloud or distributed environments [22, 9].
@@ 2 Related works
SNNs stand out from conventional neural networks by processing information through sparse, event-driven spikes rather than continuous activations. Temporal coding enables efficient handling of time-dependent data and reduces energy consumption, as neurons fire only in response to stimuli [16], [17]. The combination of energy efficiency and temporal sensitivity makes SNNs well suited for resource-constrained environments and real-time applications such as robotics, autonomous driving, and decision-making in dynamic contexts. Compared with traditional Artificial Neural Networks (ANNs), which use dense, continuous activations [34, 23], SNNs exploit spike timing and frequency to capture richer spatio-temporal patterns [36]. Such spatio-temporal encoding improves efficiency in domains including vision, speech, and medical diagnostics, while also supporting sustainable, low-power computing [56]. The temporal precision inherent in spike-based computation is particularly advantageous for secure data processing, where sparsity and accurate spike timing can enhance encryption and strengthen resilience against attacks.
@@
Synaptic plasticity modulates strength through Long Term Potentiation (LTP) and Long Term Depression (LTD), with transient dynamics captured by Short Term Potentiation (STP) and Short Term Depression (STD). A key mechanism, STDP, adjusts synapses based on pre- and post-synaptic spike timing as illustrated in Figure 3 [2, 51, 48], typically modelled with polynomial or exponential timing-dependent windows. Weight initialisation strongly affects network stability, convergence, and learning performance. Methods include random, interval-based, variance-scaling, data-driven, and hybrid approaches [45]. Balanced initialisation supports STDP, avoiding under- or over-spiking, while lateral inhibition mitigates excessive firing by suppressing neighbouring neurons. Remote Supervised Method (ReSuMe) learning uses an instructive signal to modulate plasticity without altering the post-synaptic membrane [3], ensuring precise spike timing. Effective for both single- and multi-layer networks, it surpasses gradient-based methods like SpikeProp [40] and remains important for supervised SNN tasks.
@@
The Internet of Things (IoT) integrates physical and digital systems, enabling autonomous device communication via Wireless multimedia sensor networks (WMSN), radio frequency identification (RFID), and cloud computing [47, 37]. Wireless sensor networkss (WSNs) provide real-time scalar, image, audio, and video data to support intelligent decision-making and automation. The inherently open and resource-constrained nature of IoT networks exposes them to eavesdropping, data tampering, and unauthorised access. To address these challenges, cryptography ensures data confidentiality and integrity by transforming plaintext into ciphertext using encryption keys and restoring it upon reception [39, 8].
Cryptanalysis identifies weaknesses in these algorithms, driving the advancement of cryptology and reinforcing the trustworthiness of IoT communications. Cryptanalysis identifies vulnerabilities, forming the field of cryptology [12, 27]. Symmetric cryptography employs a single shared key, requiring secret exchange, and offers high speed and low resource use (AES, DES, 3DES, Blowfish). As summarised in Table 1 Asymmetric cryptography uses public-private key pairs, supporting digital signatures and secure key exchange (RSA, Diffie-Hellman, Elliptic Curve Cryptography (ECC), Digital Signature Algorithm (DSA)). Lightweight, bio-inspired, and chaos-based schemes optimise security for constrained devices. AES processes 128-bit blocks with 10, 12, or 14 rounds depending on key size [7], using substitution, shift rows, mix columns, and add round key operations [42, 14]. RSA, introduced in 1977, relies on the difficulty of factorising large numbers and supports one-way encryption for secure transmission and digital signatures [21, 43]. Recent variants enhance wireless sensor network efficiency and biometric encryption, achieving high genuine rates with minimal false positives [10, 41]. A hybrid SNN-RSA approach enhances security and adaptability for data transmission. Input is first encrypted with RSA, processed by an SNN, and transmitted, with decryption reversing the sequence.
3 Methodology
Developing secure and efficient encryption systems is a critical challenge in cybersecurity. This article proposed the use of SNNs for bio-inspired encryption by integrating traditional methods like DES, AES, and RSA to enhance both security and performance, as illustrated in Fig. 2, the architecture draws inspiration from biological neurons, where dendrites receive input spikes, the axon propagates these electrical signals, and synapses modulate transmission strength.The NEST simulator [18], chosen for its biological accuracy, was used to simulate SNNs, while Python and libraries such as NumPy, Matplotlib, and JSON provided a robust development environment. System configurations included neuron types, simulation time steps, and synaptic plasticity models, with OpenCV used to visualise neural activity and refine the model. Data preprocessing involved converting categorical data into ASCII values suitable for SNN input. For instance, the word ”FIL” was mapped to ASCII values 77, 97, and 110, forming the basis for spike train generation and neuron input currents. Binary representations of ASCII values were transformed into spike trains, where ’1’ indicated a spike and ’0’ indicated no spike (e.g., ASCII 77 = ’01001101’). These spike trains enabled the SNN to process data in a biologically inspired manner. ASCII values were applied as input currents to influence neuronal firing rates and patterns, encoding data into distinct neural activity patterns for different characters and inputs.
@@ 3.2.7 Evaluation metrics
These metrics collectively capture both the overall accuracy and the balance between false alarms FP and missed detections FN. Precision and Recall quantify the trade-off between false positives and false negatives, while the F1-score provides a single measure that balances them. The WCR and CCR highlight overall classification reliability, which is particularly relevant when evaluating noisy spike-train data processed by the SNN.
@@ 3.3 Network Training and Testing
In the project, employ the ReSuMe algorithm [57] for training the SNN. ReSuMe is particularly well-suited for SNNs because it focuses on adjusting synaptic weights based on the precise timing of spikes, enabling the network to learn temporal patterns with high accuracy. The training involves fine-tuning the learning parameters to achieve optimal performance. For instance, the learning rate, which controls how much the synaptic weights are adjusted in response to each error. Another key consideration during training is the role of STDP, a mechanism where the timing of spikes directly influences synaptic strength. The network’s ability to adjust and refine its response to temporal patterns during training. By monitoring and adjusting STDPs effects, ensure that the network is learning effectively and not just memorising the input patterns. Once training is completed, the network undergoes a rigorous testing phase using a separate validation dataset. It is designed to evaluate how well the network has generalised its learning to new, unseen data. The performance of the SNN is measured using specific metrics such as the Purpura Distance and the Van Rossum metrics.
@@ 3.3.1 Purpura Distance
To compare the dissimilarities between spike trains from different neurons, this method introduces a concept of distance between pairs of spike trains, offering an alternative to the commonly employed rate-based correlation techniques for analysing neuronal responses. Here, spike trains of finite length are represented as points within an abstract space, where a specialised metric assigns a non-negative value $D_{ij}$ to each pair of points $i$ and $j$ [20, 58]. The Victor-Purpura distance possesses key characteristics of a true metric: it is zero only for identical spike trains ($D_{ii}=0$), positive for different spike trains ($D_{ij}>0$ for $i\neq j$), symmetric ($D_{ij}=D_{ji}$), and adheres to the triangle inequality ($D_{ik}\leq D_{ij}+D_{jk}$). The Purpura distance is calculated by determining the minimum cost required to convert one spike train into another through operations such as adding or removing spikes, shifting spike times, or changing the neuron’s identity. The cost associated with each modification is determined by the parameter $q$, which controls the timescale sensitivity for spike shifts [19]. The metric can be generalised into a family of distances, each sensitive to the neuron from which each spike originates. In the analysis, employ the basic Purpura metric with a cost of $q=1$ per unit time for shifting a spike.
@@ 4.1 Benchmark using the cryptography and PyCryptodome libraries
The adoption and development of cryptographic libraries in Python have undergone significant evolution, with libraries like PyCrypto [^1] Available online, https://www.pycrypto.org/, last accessed: 25/08/2025 and Cryptography [^2] Available online, https://github.com/pyca/cryptography, last accessed: 25/08/2025 serving as pivotal milestones. PyCrypto, one of the earliest libraries, provided developers with essential cryptographic tools, including support for symmetric encryption algorithms (e.g. AES, DES), hashing algorithms, and public-key cryptography. However, despite its foundational role, PyCrypto suffered from several critical shortcomings. The library’s reliance on insecure defaults, such as AES in Electronic Codebook (ECB) mode, and its auto-generated, poorly structured documentation left developers prone to implementation errors. These gaps not only undermined security but also forced developers to rely on external, often unreliable, sources like blogs and forums, further propagating insecure practices